OpenStudy (mikeymaximum):

Replace ? with =, >, or < to make the statement true. 2 ∙ 18 – 18 ? 2 ∙ [28 – 3 ∙ (2 + 4)] A. = B. < C. >

OpenStudy (mikeymaximum):

I got C.

OpenStudy (anonymous):

Let's start on the left side. 2*18 = 36 36-18 = 18. So we know the left side is 18.

OpenStudy (anonymous):

The right side we have the (2+4) in the parentheses so we solve that first which equals 6. now we take -3 and multiply that by 6 and that equals -18. so on the right we have 2*[28 - 18] 2*[10] = 20 The right side will equal 20. 18 < 20 Which is C. You are correct.